Disk-directed I/O for MIMD Multiprocessors

David Kotz

Department of Computer Science Dartmouth College, Hanover NH

Abstract

Many scientific applications that run on today’s multiprocessors, such as weather forecasting and seismic analysis, are bottlenecked by their file-I/O needs. Even if the multiprocessor is configured with sufficient I/O hardware, the file-system software often fails to provide the available bandwidth to the application. Although libraries and enhanced filesystem interfaces can make a significant improvement, we believe that fundamental changes are needed in the file-server software. We propose a new technique, disk-directed I/O, to allow the disk servers to determine the flow of data for maximum performance. Our simulations show that tremendous performance gains are possible. Indeed, disk-directed I/O provided consistent high performance that was largely independent of data distribution, obtained up to 93% of peak disk bandwidth, and was as much as 18 times faster than traditional parallel file systems.

Refreshments and cookies will be served